Sol Summary – April 10th

Crew 192 Sol 2 Summary Report

Sol 2

Summary Title: A Beautiful Day for EVAs

Author: Victoria Varone

Mission Status: It was a beautiful day for EVAs over here for Crew 192 on Mars. We continue to clean and organize things around the hab and have started getting more comfortable with EVAs, which will allow us to continue the search for fascinating natural wonders as well as our missing science experiments and crew member. The mission is going well.

Sol Activity Summary: The day started with a morning EVA, where three of our crew checked out the Marble Ritual site and surrounding area. They discovered some interesting geological structures and were particularly fascinated by a view of snow-capped mountains from a dusty, sandy red desert. After exploring, the crew returned to the hab, examined one of the hab’s water tanks for future cleaning, and headed inside for lunch.

After a great afternoon session of debriefing, brainstorming, and improvising for additional science experiments, the remaining crew members headed out for another EVA, coming back with excited descriptions of their hike in Kissing Camels Ridge. They also had some insight into the length of time of EVAs, noting that the amount of hiking they did in their 2.5 hour EVA expended just enough of their energy without causing serious dehydration. They’ve advise the rest of us to limit high-impact EVAs to about two and a half hours for safety.

We’re now resting and planning for tomorrow’s activities and may end the night with a space-based movie.

Look Ahead Plan: The goal is continue stretching our EVAs a little further as we keep searching for our missing science experiments and crew member. Based on tonight’s hiking EVA, we also plan to pay special attention to crew health on EVAs, making sure everyone is healthy and comfortable and no one becomes dehydrated and ill. A little attention is also being put towards possible activities for our Yuri’s Night celebration.

Anomalies in Work: None

Weather: The weather was beautiful again today, the skies started out clear and the weather cool, which worked well for the first EVA. It became overcast from the afternoon on, which allowed our second EVA crew members to explore the sites with different skies overhead.

Crew Physical Status: All crew are physically healthy and safe.

EVA: See EVA Report.

Reports to be filed: Sol Summary Report, GreenHab Report, EVA Request, Journalist Report, Commander’s Report, Photo Report

Support Requested: None

Copyright © The Mars Society. All rights reserved. | Main Site